A certain metal has a coefficient of linear expansion of 2.00 × 10-5 K-1. It has been kept in a laboratory oven at 325°C for a long time. It is now removed from the oven and placed in a freezer at -145°C. After it has reached freezer temperature, the percent change in its density during this process is closest to

A) +2.90%.

B) -2.90%.

C) +2.74%.

D) -2.74%.

The anwser is A. I need an explanation
Answer #1

ろ d d d o 木2-5-/-
Answer #2

if \alpha is the coefficient of linear expansion ,\beta is the coefficient of surface expansion and \gamma is the coefficient of volumetric expansion then the relation between the \alpha ,\beta ,\gamma is

  \alpha /1 =\beta/2 = \gamma /3 from this the relation between the \alpha ,\gamma is

  \gamma = 3\alpha   -----------------------(1)

if V1 ,d1 are the intial volume and density and V2 ,d2 are the final volume and density then

from density (d) = mass (m) / volume (V)   

d V = constant then

d1 V1 =d2 V2

d2 /d1 = V1/V2 -----------------------(2)

from the volumetric compression the final volume V2 of the metal is obtained from below equation

V2 = V1 (1 - \gamma\DeltaT) here \gamma is replace with 3\alpha from (1) we get

   V2 = V1 (1 - 3\alpha  \DeltaT) and

V2 /V1 = 1 - 3\alpha  \DeltaT   

reverse the above equation we get

V1 /V2 = 1 / (1 - 3\alpha  \DeltaT )

using (2) replace  V1 /V2 by d2/d1 in above equation then

  d2/d1 = 1 / (1 - 3\alpha  \DeltaT )

now subtract 1 from both sides of the equation we get

   (d2/d1) -1 = [1 / (1 - 3\alpha  \DeltaT ) ] - 1

(d2-d1) / d1 =(3\alpha\DeltaT) / (1 - 3\alpha  \DeltaT ) ......... { d2-d1 =\Delta d }

  \Delta d /d = (3\alpha\DeltaT) / (1 - 3\alpha  \DeltaT )

here \alpha =2.00 × 10-5 K-1   and \DeltaT = (325 - (-145 )) K =470 K (difference will be same in both celsius and kelvin)

On substituting above values we get the percent of density increased with postive value ( Here d2> d1 , because during the compression of metal volume decreases and density increases)

  (\Delta d /d ) x100 = (3 x2.00 × 10-5 K-1 x470K ) ( 1- ( 3 x2.00 × 10-5 K-1 ​x470K)) x100

  (\Delta d /d ) x100 = + 2.90 %

Therefore the % of change in density is   + 2.90 %

so option (A) is the correct one .